OpenStudy (anonymous):

Sara wants to solve 73 × 4 = . Which strategy could Sara use? A. (7 × 4) + (7 × 3) = B. (70 × 4) + (3 × 4) = C. (70 × 3) + (70 × 4) = D. (7 × 3) + (3 × 4) =

OpenStudy (anonymous):

this may not be hard for some people but i have trouble learning thing's @yrelhan4

OpenStudy (anonymous):

@jhawes

OpenStudy (yrelhan4):

Let's see.. I want to write 73 in terms of 70 and 3.. How can i write that?

OpenStudy (anonymous):

like i said trouble with math im not getting it

OpenStudy (anonymous):

i have a learning disability

OpenStudy (yrelhan4):

Forget the original problem. What operator should I use between 70 and 3 to make 73.. 70 ? 3 Options: +(addition) -(subtraction) *(multiplication) /(division)

OpenStudy (anonymous):

+

OpenStudy (anonymous):

B because 73*4=(70+3)*4=(70*4)+(3*4)

OpenStudy (yrelhan4):

Good. so the original equation becomes (70+3)*4.. following me?

OpenStudy (yrelhan4):

You're not supposed to give direct answers. Defeats the purpose.

OpenStudy (anonymous):

@minniepotato I would say B you just split 73 into two things so 70 and 3 so it would be 70 x 4 and 3x4

OpenStudy (anonymous):

thank you guys for explaining it also